# Is Your LLM Outdated? Evaluating LLMs at Temporal Generalization

### v2

@@ Appendix A Testing Statistical Hypothesis
Due to the inherent randomness in data, directly comparing outcomes of two experiments is insufficient to justify the superiority of one outcome over the other. To address this issue, practitioners often resort to hypothesis testing to make convincing decisions, like those demonstrated in [27]. A board application of hypothesis testing lies in the A/B testing literature, which assesses the performance of various features of a product and performs post-experiment inferences, as introduced in [28].
Appendix B Details of Data Collection
@@
Appendix C Data Overview of One Crawl
@@
Category Source Count Length
Academic STEM arXiv Mathematics 50 42850
arXiv Computer Science 50 43336
arXiv Physics 50 42580
arXiv Statistics 50 42619
arXiv Electrical Engineering and Systems Science 50 43027
Academic Non-STEM arXiv Economics 50 42592
arXiv Quantitative Finance 50 43969
arXiv Quantitative Biology 50 42974
Code GitHub 30 21849
Internet QA Reddit 583 643
Quora Health 20 1150
Quora Movies 69 1349
Quora Mathematics 70 1365
Quora Technology 80 1674
Internet article Wiki 220 4091
Wattpad 38 6615
BBC 45 4792
Yahoo 41 5373
Appendix D Detailed results of BPC generalization on Wiki and BBC dataset
@@
Appendix E TBI of Models across the arXiv, BBC, and Wiki datasets
@@
Appendix F Most common tags of the Gjo questions
Table 8: Frequency of Topics
Topic Count
Business 648
Non-US Politics 453
Security and Conflict 446
Society 425
US Politics 363
Technology 359
Finance 299
Elections and Referenda 288
Economic Indicators 273
Foreign Policy 271
US Policy 247
Health 245
Leader Entry/Exit 219
Economic Policy 160
Sports 148
Entertainment 148
Environment 80
Open 16
Table 9: Most common tags of the Gjo questions
Appendix G Correlation with Existing Benchmarks
Following our previous chapter I, we continue to explore the correlation between prediction performance and existing benchmarks, using the same benchmarks as in I. We have categorized the tags of questions into several groups: Societal Issues Items, Political and Government Affairs, Economic and Business Items, and Science and Technology Items. Each group represents a significant area of interest.
In our analysis, we focus on how the correlation between these categories and various benchmarks evolves over time. We use correlation coefficients to measure the strength of association between the performance scores of questions within each category and the benchmarks across different times.
Notably, our analysis revealed that while the Economic and Business Items and Science and Technology Items on base models displayed relatively high initial correlations, there appears to be little to no relationship in Societal issues Items and Political and Government Affairs, we think such topics are influenced by a multitude of unpredictable factors such as social changes, policy adjustments, and public sentiment. These elements make accurate predictions particularly challenging, leading to the observed low correlation with performance metrics. Detailed figures are in Appendix N.
Despite relatively high initial correlations in 2 question groups, there is a general lack of correlation between model predictions and existing benchmarks (mean correlation coefficient around zero). In addition, significant deviations occur in nearly half of the cases from early stages to subsequent periods. This suggests that the models’ understanding and alignment with benchmarks can vary substantially over time. So temporal generalization provides novel insights and draws attention to an aspect that has been often ignored by previous benchmarks.
Appendix H Analysis of Language Likelihood Values and Benchmark Correlation Across Text and Model Groups
@@
Appendix I Correlation with Existing Benchmarks
@@
HellaSwag Tests an AI’s common sense reasoning by requiring it to complete sentences or narratives that reflect everyday logic, highlighting the model’s ability to predict logical continuations. [29]
MMLU Evaluates an AI’s breadth and depth of understanding across a diverse range of subjects, from science to literature, showcasing its versatility and comprehensive grasp of human language. [30]
TruthfulQA Focuses on the model’s capacity to provide honest and accurate responses, particularly to questions where the truth may not be intuitive, testing its commitment to truthfulness and its proficiency in avoiding misinformation. [31]
Winogrande Assesses common sense reasoning and natural language understanding through sentence completion tasks, measuring an AI’s ability to apply common sense knowledge in language processing. [32]
GSM8K Challenges AI models with elementary-level math problems, testing their mathematical reasoning capabilities and their understanding and application of basic math concepts. [33]
@@
On the Genres of Corpora
As can be seen in Appendix  H, whether all models are observed or only those of a specific scale does not make a big difference. Comparisons across different model sizes show that the differences are not substantial, indicating consistency in performance correlations across benchmarks.
Further analysis from Tab. 13 within the MMLU benchmark sections reveals interesting patterns. While we assume models to be generally excel in domains closely aligned with their training data, this correlation is not strong. Specifically, STEM scores do not show strong correlation with STEM-related texts. Moreover, academic papers, even in Social Sciences and Humanities, exhibit stronger correlations than Internet data, which includes forums and news.
In this context, code-related data show similarities in performance correlation with academic texts, suggesting a closer alignment than previously anticipated. Similarly, Internet data, including forums and news, also shows comparative closeness in performance, indicating nuanced relationships across different types of content.
Table 13: Correlation Coefficients between BPC in specific areas and Domain-Specific Performance
Benchmark Academia STEM Academia non-STEM Code Internet Article Internet Chat
MMLU:STEM -0.17 -0.19 -0.11 -0.01 -0.01
MMLU:Humanities -0.54 -0.53 -0.51 -0.29 -0.39
MMLU:Social Sciences -0.49 -0.42 -0.46 -0.25 -0.31
Additionally, compared to chat forum texts, more formal academic texts demonstrate a stronger correlation with benchmarks. The genre of text has a minor impact on the sign of the correlation coefficient, which is relatively small compared to the differences across model types.
Appendix J Function Fitting Strategies
We considered several fitting strategies, each with its unique mathematical approach to model decay:
Linear Decay
Linear decay is represented by a straight line, suggesting a constant rate of decay over time. The model is given by the equation $y=ax+b$.
Exponential Decay
Exponential decay is modeled by an exponential function, expressed as $y=y_{0}\cdot e^{-kt}$, where $y_{0}$ is the initial value, $t$ is time or another decaying factor, and $k$ is the decay constant that denotes the rate of decay.
Power Law Decay
Power law decay implies a polynomial relationship and is expressed as $y=at^{-k}$, where $a$ is a positive constant, $t$ is time or another related factor, and $k$ represents the exponent characterizing the rate of decay.
Logarithmic Decay
Logarithmic decay suggests a decrease that becomes more gradual over time, typically modeled as $y=a+b\log(x)$, where $a$ and $b$ are constants, and $x$ represents time or another related factor.
The variance of the residuals for each model was calculated as follows:
Table 14: Variance of Residuals for Different Models
Model Variance
Linear 2.346e-3
Exponential 1.71e3
Power Law 3.78e-3
Logarithmic 3.33e-3
J.0.1 Findings
Our analysis found that the linear model provided the best fit, with the lowest variance among the residuals. This indicates a steady, unchanging rate of knowledge decay over time, as opposed to the more complex models that assume varying rates.
We aimed to evaluate the understanding and recency of knowledge across various language models, including GPT-4 and others. To accomplish this, we devised a methodology to calculate a weighted accuracy score for each model, reflecting not only its performance but also the contemporaneity of its knowledge base.
We began by assigning weights to different quarters, starting from 0 for the first quarter of 2015, up to 1 for the most recent quarter. This linear interpolation of weights was chosen to emphasize the value of more current information, underlining our hypothesis that accuracy concerning more recent knowledge is indicative of a model’s modernity and relevance.
The formula applied was straightforward: for each model, we multiplied its accuracy (Acc) by the corresponding weight based on the time difference since the first quarter of 2015. The resulting weighted accuracies were then summed up to yield an overall score for each model, which we used to rank them accordingly.
GPT-4 and other leading models demonstrated a notably later peak in knowledge, suggesting that their databases include more recent information. This is particularly valuable in fast-evolving fields where the timeliness of information is crucial.
Below is a partial presentation of rankings and scores for selected models, showcasing their weighted accuracy scores. The complete table can be found in Table 15.
Table 15: Model Scores
Model Name Score
GPT4-1106 794.275
GPT4-0613 699.8
Zephyr-7B-beta 601.35
InternLM-Chat-7B 592.15
GPT3.5-turbo-0613 578.025
Baichuan-13B-Chat 571.8
Qwen-14B-Chat 565.45
OPT-2.7B 547.375
Yi-6B-Chat 541.05
Mistral-7B-v0.1 528.725
Baichuan-7B-Chat 527.825
LLaMA-7B 524.075
Baichuan2-7B-Chat 517.725
Pythia-12B 513.375
Colossal-LLaMA2-7B-Base 513.35
LLaMA2-7B-Chat 508.4
LLaMA2-13B 508.15
Qwen-7B 506.7
Qwen-1.8B-Chat 505.225
Phi-2 497.4
Yi-6B 496.15
Zhongjing-Base 490.1
LLaMA2-7B 486.2
Falcon-rw-1B 481.15
ChatGLM3-6B 479.8
Baichuan2-7B-Base 479.175
Baichuan2-13B-Chat 463.75
OPT-13B 453.075
Baichuan2-13B-Base 446.225
Skywork-13B-Base 445.8
Qwen-1 8B 435.775
RWKV5-Eagle-7B 408.425
Qwen-7B-Chat 394.9
Phi-1.5 385.975
TinyLLaMA-1.1B-Chat-v0.6 366.125
Appendix K Details of Model Accuracy Peaks
@@
Appendix L Question Tag Details
Table 17: Detailed breakdown and sum of entries per category within each group of benchmarks.
Group Category Number of Entries
Political and Government Affairs Leader Entry/Exit 199
US Politics 338
Elections and Referenda 265
US Policy 223
Foreign Policy 242
Non-US Politics 415
Security and Conflict 413
Total: 2095
Economic and Business Items Business 575
Economic Indicators 253
Finance 284
Economic Policy 147
Total: 1259
Societal issues Items Society 379
Health 237
Environment 75
Sports 120
Entertainment 132
Total: 943
Science and Technology Items Technology 346
Total: 346
Appendix M Analysis of Best-Achieved Accuracy and Cutoff Time
@@
Appendix N Prediction Correlation over Time
The graphical representations in our analysis plot the changes in correlation coefficients over time. Each line in the graph represents the changing relationship between a category’s average adjusted performance and a specific benchmark, highlighting periods of stronger or weaker association.
Figure 7: Science and Technology Items - All Model Spearman Smoothed
Figure 8: Science and Technology Items - Base Model Spearman Smoothed
Figure 9: Societal Issues Items - All Model Spearman Smoothed
Figure 10: Political and Government Affairs - All Model Spearman Smoothed
Figure 11: Economic and Business Items - All Model Spearman Smoothed

### v3

@@ Appendix E Testing Statistical Hypothesis
Due to the inherent randomness in data, directly comparing outcomes of two experiments is insufficient to justify the superiority of one outcome over the other. To address this issue, practitioners often resort to hypothesis testing to make convincing decisions, like those demonstrated in Zhou et al. (2016). A board application of hypothesis testing lies in the A/B testing literature, which assesses the performance of various features of a product and performs post-experiment inferences, as introduced in Russac et al. (2021).
Appendix F Details of Data Collection
@@
Table 5: Most common tags of the Gjo questions
Topic Count
Business 648
Non-US Politics 453
Security and Conflict 446
Society 425
US Politics 363
Technology 359
Finance 299
Elections and Referenda 288
Economic Indicators 273
Foreign Policy 271
US Policy 247
Health 245
Leader Entry/Exit 219
Economic Policy 160
Sports 148
Entertainment 148
Environment 80
Open 16
Appendix G Data Overview of One Crawl
@@
Appendix H Detailed results of BPC generalization on Wiki and BBC dataset
See Table 6.
@@ Appendix H Detailed results of BPC generalization on Wiki and BBC dataset
Appendix I TBI of Models across the arXiv, BBC, and Wiki datasets
See Table 8.
@@ Appendix I TBI of Models across the arXiv, BBC, and Wiki datasets
Category Source Count Length
Academic STEM arXiv Mathematics 50 42850
arXiv Computer Science 50 43336
arXiv Physics 50 42580
arXiv Statistics 50 42619
arXiv Electrical Engineering and Systems Science 50 43027
Academic Non-STEM arXiv Economics 50 42592
arXiv Quantitative Finance 50 43969
arXiv Quantitative Biology 50 42974
Code GitHub 30 21849
Internet QA Reddit 583 643
Quora Health 20 1150
Quora Movies 69 1349
Quora Mathematics 70 1365
Quora Technology 80 1674
Internet article Wiki 220 4091
Wattpad 38 6615
BBC 45 4792
Yahoo 41 5373
Appendix J Gjo data samples and most common tags of the Gjo questions
Gjo data samples see Table 5 and Gjo questions see Table 10.
Table 10: Sample of Questions and Their Respective Answers (Correct Answer in Bold): The percentages represent the proportion of forecasters’ votes on a prediction website.
Question Start Time End Time Possible Answers
What will be the 12-month percentage change in the US Consumer Price Index (CPI) for September 2024? May 31, 2024 05:00PM UTC Oct 01, 2024 07:01AM UTC Up by less than 1.200% or down: 0% Up by at least 1.200%, but less than 1.800%: 0% Up by at least 1.800%, but less than 2.400%: 53% Up by at least 2.400%, but less than 3.000%: 44% Up by at least 3.000%, but less than 3.600%: 3% Up at least 3.600%, but less than 4.200%: 0% Up by 4.200% or more: 0%
If Israeli Defense Forces (IDF) ground forces invade the Gaza Strip before 7 November 2023, when will Israel publicly announce or acknowledge that IDF ground forces have left the Gaza Strip? Oct 17, 2023 05:00PM UTC Oct 10, 2024 07:01AM UTC Before 30 November 2023: 0% Between 30 November 2023 and 13 January 2024: 0% Between 14 January 2024 and 13 March 2024: 0% Between 14 March 2024 and 11 June 2024: 0% Between 12 June 2024 and 9 October 2024: 2% Not before 10 October 2024: 96% IDF ground forces will not invade the Gaza Strip before 7 November 2023: 2%
What will Kamala Harris’ favorability rating be as of 3 October 2024, according to FiveThirtyEight? Sep 17, 2024 08:00AM UTC Oct 03, 2024 07:01AM UTC Lower than 42.0%: 1% At least 42.0%, but less than 44.0%: 4% At least 44.0%, but less than 46.0%: 7% At least 46.0%, but less than 48.0%: 70% At least 48.0%, but less than 50.0%: 15% At least 50.0%, but less than 52.0%: 2% At least 52.0%, but less than 54.0%: 1% 54.0% or higher: 0%
Appendix K Analysis of Language Likelihood Values and Benchmark Correlation Across Text and Model Groups
@@
Appendix L Correlation with Existing Benchmarks
@@
HellaSwag Tests an AI’s common sense reasoning by requiring it to complete sentences or narratives that reflect everyday logic, highlighting the model’s ability to predict logical continuations. Zellers et al. (2019)
MMLU Evaluates an AI’s breadth and depth of understanding across a diverse range of subjects, from science to literature, showcasing its versatility and comprehensive grasp of human language. Hendrycks et al. (2021)
TruthfulQA Focuses on the model’s capacity to provide honest and accurate responses, particularly to questions where the truth may not be intuitive, testing its commitment to truthfulness and its proficiency in avoiding misinformation. Lin et al. (2022)
Winogrande Assesses common sense reasoning and natural language understanding through sentence completion tasks, measuring an AI’s ability to apply common sense knowledge in language processing. Sakaguchi et al. (2019)
GSM8K Challenges AI models with elementary-level math problems, testing their mathematical reasoning capabilities and their understanding and application of basic math concepts. Cobbe et al. (2021)
@@
Appendix M Details of Model Accuracy Peaks
@@
Appendix N Analysis of Best-Achieved Accuracy and Cutoff Time